e Water enters a constant 100 [mon ] I.D. tubes et a bilen at 7 [MPa] and 60000] and leaves the tubes at 6[MPa ] and 400 [°C please exlain how are you finding the specific volumes for each one to use

VA/v = (VA/v)
Answer #1

Viven P,=YMP2 Zh=251.04 kilkg Ti = 60°C Jui= P (MPa) u (m3/kg) T:+60°C 5 0.0010149 7 V 10 0.0010124 (Table A-7) 7-5 = U-0.001mass flow rate, mie PA, V, = B A2 V2 AINI - Ave